In any classroom, there is a 92% chance that a randomly selected student will ask to go to the bathroom or go to the nurse. If the probability that he or she will ask to go to the bathroom is 80%, what is the probability that he or she will ask to go to the nurse? Assume independence.
Answer by venugopalramana(3286)

IF THE EVENTS ARE INDEPENDENT THEN
P(A) U P(B)=P(A)+P(B)
WE HAVE
In any classroom, there is a 92% chance that a randomly selected student will ask to go to the bathroom or go to the nurse
FROM WHAT IS DESIGNATED BELOW THIS MEANS P(A) U P(B)=92%=0.92
. If the probability that he or she will ask to go to the bathroom
CALL IT EVENT A
is 80%,
P(A)=80/100=0.8
what is the probability that he or she will ask to go to the nurse?
CALL THIS EVENT B AND P(B)=?
Assume independence.
SO
0.92=0.8+P(B)
P(B)=0.12=12%